Jefferson Lab HALL A Logbook for May 06 - All EXCEPT Automatic Entries

  • Main INDEX

    Friday

  • 173772 05/26/06 15:18 R. Michaels duplicate GeN file cleanup
  • Thursday

  • 173771 05/18/06 16:03 gilman fpp cosmic run 4676
  • 173770 05/18/06 14:28 gilman end FPP cosmic run 4675
  • 173769 05/18/06 10:52 gilman FPP HV 8.1.12
  • 173768 05/18/06 10:48 gilman FPP cosmic test run 4675
  • Tuesday

  • 173767 05/16/06 15:32 gilman FPP cosmic test run
  • 173766 05/16/06 13:55 gilman cosmic run - FPP test
  • Friday

  • 173765 05/12/06 16:47 atrig Trigger-Download
  • 173764 05/12/06 08:09 E.Chudakov Moller/calorimeter tests
  • 173762 05/12/06 04:46 E.Chudakov Start Moller setting
  • 173761 05/12/06 04:16 B. Craver Power down detectors
  • 173758 05/12/06 04:11 sawatzky Shift/Run summary (00:00 - 04:00)
  • 173756 05/12/06 04:01 hyao Target Moving List
  • 173753 05/12/06 03:52 B. Craver MWDC currents
  • 173752 05/12/06 03:51 sawatzky Rates for 6 uA on Hydrogen cell (136 psi)
  • 173747 05/12/06 03:27 B. Craver MWDC currents
  • 173746 05/12/06 03:24 sawatzky Rates for 4 uA on Hydrogen
  • 173736 05/12/06 02:18 sawatzky Run 4650 is 1uA on Hydrogen cell (136 psi)
  • 173735 05/12/06 02:18 sawatzky Run 4649 is 1 uA Hydrogen cell (136 psi)
  • 173734 05/12/06 02:17 sawatzky Run 4648 is 1uA optics foil
  • 173733 05/12/06 02:02 sawatzky Rates for 1uA on H targ (136 psi)
  • 173728 05/12/06 01:45 Adaq Raster run 4648
  • 173717 05/12/06 00:19 X. Jiang swing shift summary. test run
  • Thursday

  • 173701 05/11/06 22:37 Xin Qian Momentum reconstruction
  • 173697 05/11/06 21:48 A. Camsonne HV genps1 2001 crashed
  • 173690 05/11/06 21:08 Xin Qian naive carbon foil reconstruction
  • 173683 05/11/06 20:11 Xin Qian raster pattern
  • 173679 05/11/06 20:04 Xin Qian Run 4634
  • 173675 05/11/06 19:56 Xin Qian raster pattern
  • 173667 05/11/06 19:23 E.Chudakov GeN surveys
  • 173664 05/11/06 19:05 Xin Qian Turn On power supply for pre-amplify card.
  • 173660 05/11/06 18:46 x. jiang neutron arm to 32.5
  • 173608 05/11/06 01:22 A. Kolarkar Target cold; lasers locked out
  • 173607 05/11/06 01:20 JP, Ameya, X the laser power measurement is done
  • Wednesday

  • 173600 05/10/06 23:16 A. Kolarkar Target Studies
  • 173583 05/10/06 12:09 Katich lasers off, oven to 0, etc...
  • 173581 05/10/06 07:25 kelleher target to 3He position
  • 173580 05/10/06 07:25 kelleher Big Bite ON
  • 173579 05/10/06 07:24 E.Chudakov Moller/calorimeter tests
  • 173578 05/10/06 00:39 frw target on empty
  • 173577 05/10/06 00:26 X. Jiang BigBite off.
  • 173576 05/10/06 00:26 x. jiang run 4598, good optics foil run.
  • 173573 05/10/06 00:14 X. Jiang run 4598, wirechamber HV screen
  • 173571 05/10/06 00:03 frw target to optics foils
  • Tuesday

  • 173570 05/09/06 23:57 S.ABrahamyan Shift Summary 16-24
  • 173567 05/09/06 23:30 x. jiang spot++ during run 4597
  • 173563 05/09/06 23:16 A. Kolarkar NMR measurement: Pol = 53.5 %
  • 173562 05/09/06 23:12 A. Tobias Moved to H2 Reference Cell (135psig)
  • 173558 05/09/06 22:46 S.Abrahamyan BHWP IN
  • 173548 05/09/06 19:48 A. Tobias NMR measurement: Pol = 55.2 %
  • 173544 05/09/06 19:02 A. Tobias Upper Lockin not on during an NMR @ 17:55
  • 173533 05/09/06 16:04 B. Craver Day shift summary
  • 173529 05/09/06 14:35 B. Craver MWDC currents
  • 173528 05/09/06 14:33 B. Craver Return to production - spot++ screen
  • 173524 05/09/06 14:27 B. Craver MWDC currents
  • 173523 05/09/06 14:26 B. Craver spot++ screen shot
  • 173518 05/09/06 14:20 B. Craver MWDC currents
  • 173517 05/09/06 14:20 B. Craver spot++ screen shot
  • 173516 05/09/06 14:16 Katich target moved
  • 173513 05/09/06 13:32 B. Craver MWDC Currents
  • 173511 05/09/06 13:27 B. Craver Reference cell pressure
  • 173504 05/09/06 13:09 B. Craver MWDC HV screen shot
  • 173503 05/09/06 13:08 Katich target moved to reference
  • 173500 05/09/06 12:28 tajima analysis shift summary 8-12 (FINAL!!!)
  • 173498 05/09/06 12:26 B. Craver MWDC currents
  • 173497 05/09/06 12:25 Katich NMR measurement: Pol = 55.4 %
  • 173493 05/09/06 12:14 B. Craver optics data
  • 173492 05/09/06 12:02 B. Craver beam down
  • 173486 05/09/06 10:19 tajima Strange Qasym result for run 4576
  • 173482 05/09/06 09:18 B. Craver Fine tune MWDC voltages
  • 173478 05/09/06 09:03 A. Camsonne Analysis shift 6-9 AM
  • 173474 05/09/06 08:34 S. Frullani owl shift summary
  • 173472 05/09/06 07:54 S. Frullani BWHP OUT
  • 173462 05/09/06 06:02 J. Miller Analysis Shift 2 am - 6 am
  • 173460 05/09/06 05:50 frw NMR measurement: Pol = 55.4 %
  • 173454 05/09/06 03:33 A. Kolarkar Analysis Shift Summary 2300 - 0300
  • 173441 05/09/06 00:18 Nerses Shift summary 20:00-24:00
  • 173440 05/09/06 00:02 B. Crowe Swing Shift Run Summary
  • Monday

  • 173435 05/08/06 23:11 B. Quinn NMR measurement: Pol = 55.7 %
  • 173428 05/08/06 21:06 S.Abrahamyan Shift 17-21
  • 173424 05/08/06 20:13 B. Crowe BPMB-Y adjustment
  • 173414 05/08/06 16:56 B. Quinn NMR measurement: Pol = 55.2 %
  • 173410 05/08/06 16:23 B. Crowe BHWP is IN
  • 173408 05/08/06 16:00 Saha Day Shift Summary
  • 173401 05/08/06 13:45 tajima analysis shift summary 8-12
  • 173391 05/08/06 10:26 protopop NMR measurement: Pol = 55.3 %
  • 173384 05/08/06 08:34 A. Camsonne Shift summary 6-9AM
  • 173380 05/08/06 08:02 Armando Acha Owl Shift Summary
  • 173378 05/08/06 07:47 A. Acha BTA correction
  • 173376 05/08/06 07:30 A. Acha BTA mistake
  • 173375 05/08/06 07:14 J. Miller Neutron HV Trip
  • 173369 05/08/06 05:35 J. Miller Analysis Shift 2 am - 6 am
  • 173365 05/08/06 04:31 Armando Acha Raster check. Run 4548
  • 173364 05/08/06 04:29 A. Acha MWDC Currents. Run 4548
  • 173363 05/08/06 04:26 Armando Acha Beam is back. Run 4548
  • 173358 05/08/06 03:31 A. Kolarkar Analysis Shift Summary 2300 - 0300
  • 173357 05/08/06 03:15 A. Kolarkar NMR measurement: Pol = 54.5 %
  • 173356 05/08/06 02:51 Armando Acha Run 4546 stopped. No beam
  • 173347 05/08/06 00:52 Armando Acha Happex FM snapshot
  • 173346 05/08/06 00:51 A. Acha raster size. run 4544
  • 173345 05/08/06 00:44 Nerses Shift summary 20:00-24:00
  • 173344 05/08/06 00:02 B. Crowe Swing Shift Run Summary
  • Sunday

  • 173342 05/07/06 23:40 B. Crowe Beam Restored-Raster Scan
  • 173340 05/07/06 23:22 B. Crowe No Beam/ MCC are fixing Problem with RF
  • 173334 05/07/06 22:00 B. Crowe Snapshot of Neutron detector rates for Runs 4539-40
  • 173333 05/07/06 21:38 A. Camsonne F1 TDC board stuck
  • 173332 05/07/06 21:33 S.Abrahamyan Shift 17-21
  • 173328 05/07/06 21:28 S.Abrahamyan Problem with neutron detector
  • 173327 05/07/06 21:08 B. Quinn NMR measurement: Pol = 55.3 %
  • 173326 05/07/06 20:57 B. Crowe Problems With Neutron Detector HV
  • 173319 05/07/06 19:39 B. Crowe BMPB-Y Adjustment
  • 173318 05/07/06 19:20 B. Crowe End run failed. Rebooted Coda
  • 173303 05/07/06 16:02 riordan Day Shift Summary
  • 173301 05/07/06 15:16 protopop NMR measurement: Pol = 55.1 %
  • 173299 05/07/06 15:03 riordan 4533 End of Run Failed
  • 173295 05/07/06 13:32 riordan TS1 Crashed
  • 173290 05/07/06 12:36 tajima analysis shift summary 8-12
  • 173286 05/07/06 11:03 riordan Drift Chamber Current
  • 173279 05/07/06 10:02 A. Camsonne Shift summary 6-9 AM
  • 173277 05/07/06 09:38 riordan Beam Restored - Raster++
  • 173276 05/07/06 09:36 A. Camsonne Thresholds set for data taking
  • 173275 05/07/06 09:32 kelleher NMR measurement: Pol = 55.3 %
  • 173272 05/07/06 08:44 A. Camsonne Compton lower background period
  • 173271 05/07/06 08:01 Armando Acha Owl Shift Summary
  • 173266 05/07/06 07:34 A. Camsonne Threshold changed for cosmics for neutron detector
  • 173264 05/07/06 07:17 A. Acha Not beam yet.
  • 173263 05/07/06 06:23 J. Miller Analysis Shift 2 am - 6 am
  • 173262 05/07/06 05:57 Armando Acha No beam yet
  • 173261 05/07/06 05:44 Armando Acha BHWP out.
  • 173260 05/07/06 05:04 A. Acha Run 4524 stopped. No beam.
  • 173257 05/07/06 04:08 Armando Acha Run 4523. DAQ ROC 3 crashed.
  • 173255 05/07/06 03:39 S. Frullani NMR measurement: Pol = 54.9 %
  • 173251 05/07/06 02:17 Armando Acha MWDC currents. RUN 4522
  • 173250 05/07/06 02:14 A. Acha Run 4521. ROC3 crashed
  • 173245 05/07/06 01:49 Armando Acha BCM view
  • 173244 05/07/06 01:41 Armando Acha Happex FM snapshot
  • 173243 05/07/06 01:26 Nerses Shift summary 20:00-24:00
  • 173238 05/07/06 00:01 B. Crowe Swing Shift Run Summary
  • Saturday

  • 173237 05/06/06 23:43 Nerses raster after the beam came back
  • 173236 05/06/06 23:41 Nerses Beam position
  • 173234 05/06/06 23:31 Nerses Beam trip at the end of 4518 run
  • 173233 05/06/06 23:24 B. Crowe MCC working on Fast Feed Back
  • 173227 05/06/06 22:03 B. Crowe Double peak in spectrum 11-total sum 22 in rn 4517
  • 173225 05/06/06 21:40 Nerses The RF time distribution was changed
  • 173222 05/06/06 21:02 bogdanw N-arm scaler display
  • 173220 05/06/06 20:48 B. Quinn NMR measurement: Pol = 52.9 %
  • 173211 05/06/06 18:09 B. Craver Shift summary 2-6
  • 173210 05/06/06 18:04 B. Crowe Spectrum 18-N1 from run 4512
  • 173209 05/06/06 17:53 B. Crowe Changes in replay spectrum 18-N1 and 17-V1 left
  • 173202 05/06/06 16:03 riordan Day Shift Summary
  • 173197 05/06/06 14:55 kelleher NMR measurement: Pol = 51.9 %
  • 173195 05/06/06 14:43 A. Kolarkar Target Lasers Helicity Flipped
  • 173185 05/06/06 12:42 tajima analysis shift summary 8-12
  • 173181 05/06/06 11:10 riordan Raster++ Check
  • 173177 05/06/06 10:41 riordan BPMB Y Position Corrected
  • 173176 05/06/06 10:03 riordan Half Wave Plate Moved to In
  • 173172 05/06/06 09:54 riordan Drift Chamber Currents - Run 4503
  • 173171 05/06/06 09:34 A. Camsonne Shift analysis 5-9 AM
  • 173170 05/06/06 09:09 riordan Beam Position on BPMB
  • 173166 05/06/06 08:32 riordan HAPPEX Feedback Monitor
  • 173162 05/06/06 08:03 Armando Acha Owl Shift Summary
  • 173160 05/06/06 07:17 S. Frullani NMR measurement: Pol = 53.3 %
  • 173153 05/06/06 05:42 Nerses shift summary 20:00-05:00
  • 173152 05/06/06 05:22 A. Camsonne Compton vertical scan
  • 173151 05/06/06 05:16 A. Acha Run 4498. Coda crashed
  • 173146 05/06/06 03:48 A. Acha MWDC Currents. run 4497
  • 173145 05/06/06 03:13 A. Acha Happex FM view
  • 173144 05/06/06 03:10 Armando Acha BCM snapshot
  • 173137 05/06/06 01:26 S. Frullani NMR measurement: Pol = 53.4 %
  • 173135 05/06/06 00:54 A. Acha Happex FM snapshot. HWPt OUT
  • 173132 05/06/06 00:30 A. Camsonne Compton background rates dropped
  • 173131 05/06/06 00:01 Xiaofeng Shift summary
  • Friday

  • 173126 05/05/06 23:48 Xiaofeng run 4493: compton tuning
  • 173116 05/05/06 20:17 S.Abrahamyan Shift 17-21
  • 173109 05/05/06 18:37 S.Abrahamyan Asymetry for runs 4486, 4487, 4488
  • 173108 05/05/06 18:33 Xiaofeng CODA data taking: event limit changs to 3M
  • 173107 05/05/06 18:31 B. Quinn NMR measurement: Pol = 52.9 %
  • 173103 05/05/06 17:36 Xiaofeng Beam half wave plate : in
  • 173093 05/05/06 16:26 rsholmes Half wave plate changes
  • 173092 05/05/06 16:12 rsholmes Half wave plate changes
  • 173091 05/05/06 16:04 B. Crowe Day Shift Run Summary
  • 173086 05/05/06 15:27 B. Crowe Rebooted ROC25 after crash
  • 173073 05/05/06 13:00 B. Crowe Half-Wave Plate Out
  • 173069 05/05/06 12:26 tajima analysis shift summary 8-12
  • 173068 05/05/06 12:09 protopop NMR measurement: Pol = 53.5 %
  • 173064 05/05/06 11:20 B. Crowe Rebooted Coda/ROC4 not connected
  • 173056 05/05/06 09:28 A. Camsonne Shift summary 6-9 AM
  • 173054 05/05/06 09:25 B. Crowe End Run failed
  • 173046 05/05/06 08:43 tajima ND HV trip
  • 173040 05/05/06 08:01 Armando Acha Owl Shift Summary
  • 173038 05/05/06 07:36 A. Acha Compton Polarimeter view
  • 173034 05/05/06 07:32 A. Acha Big Bite view
  • 173033 05/05/06 07:30 Armando Acha MWDC Currents. Run 4469
  • 173032 05/05/06 07:29 A. Acha Beam line information
  • 173031 05/05/06 07:28 Armando Acha BCM snapshot
  • 173030 05/05/06 07:26 A. Acha Happex FM snapshot
  • 173026 05/05/06 06:15 A. Kolarkar Analysis Shift Summary 2300 -
  • 173022 05/05/06 05:49 S. Frullani NMR measurement: Pol = 54.0 %
  • 173018 05/05/06 05:10 Armando Acha MWDC snapshot
  • 173017 05/05/06 04:40 Armando Acha Charge Asymetries. run 4465
  • 173010 05/05/06 03:30 Armando Acha Compton Polarimeter Snapshot
  • 173006 05/05/06 02:42 A. Acha Compton Polarimeter Monitor
  • 173002 05/05/06 02:03 A. Acha Compton Polarimeter Monitor
  • 173001 05/05/06 01:46 A. Acha BCM snapshot
  • 172994 05/05/06 00:54 A. Acha raster size . Run 4460
  • 172993 05/05/06 00:53 A. Acha Hapeex FM snapshot
  • 172992 05/05/06 00:37 Nerses Shift summary 20-24
  • 172988 05/05/06 00:01 Xiaofeng Shift summary
  • Thursday

  • 172980 05/04/06 22:31 B. Quinn NMR measurement: Pol = 54.1 %
  • 172976 05/04/06 21:33 S.Abrahamyan Shift 17-21
  • 172972 05/04/06 21:16 Xiaofeng After compton tuning ,data production resumes since run 4451 at 18:13
  • 172965 05/04/06 19:54 A. Kolarkar Target Polarization States for Edna
  • 172955 05/04/06 18:02 B. Quinn NMR measurement: Pol = 53.5 %
  • 172954 05/04/06 17:41 Xiaofeng beam half waveplate : in; target polarization is parallel to the magnetic field
  • 172943 05/04/06 16:00 Saha Day Shift Summary
  • 172939 05/04/06 15:13 A. Camsonne Compton background rates
  • 172938 05/04/06 14:46 S.Abrahamyan go back to 6uA
  • 172934 05/04/06 14:40 S.Abrahamyan MWDC HV
  • 172923 05/04/06 12:17 tajima analysis shift summary 8-12
  • 172922 05/04/06 11:12 S.Abrahamyan MWDC HV
  • 172918 05/04/06 10:38 A. Camsonne CODA restart and ROC reboot
  • 172917 05/04/06 10:38 gilman NMR measurement: Pol = 54.4 %
  • 172916 05/04/06 10:28 tajima ND & BB HV trips
  • 172911 05/04/06 09:23 gilman beam heats target
  • 172905 05/04/06 08:45 A. Camsonne Analysis shift summary 6-9 AM
  • 172904 05/04/06 08:43 A. Camsonne Charge asymmetry change and Half wave plate change during run 4434
  • 172901 05/04/06 08:07 Armando Acha Owl Shift Summary
  • 172900 05/04/06 08:05 A. Acha Happex feedback monitor
  • 172891 05/04/06 06:41 Armando Acha Beam Monitor
  • 172890 05/04/06 06:40 A. Acha Happex feedback monitor
  • 172889 05/04/06 06:34 J. Miller Analysis Shift Summary 2 am - 6 am
  • 172888 05/04/06 06:28 Armando Acha run 4432. production
  • 172887 05/04/06 06:21 A. Acha raster size in production mode
  • 172883 05/04/06 05:46 A. Acha MWDC Currents. Run 4431
  • 172882 05/04/06 05:45 Armando Acha raster size 1.5x1.5. run 4431
  • 172878 05/04/06 05:19 A. Acha MWDC Currents. Run 4430
  • 172877 05/04/06 05:17 Armando Acha raster size 2.5x2.5. run 4430
  • 172873 05/04/06 05:01 Armando Acha MWDC Currents. Run 4429
  • 172872 05/04/06 04:59 A. Acha Raster 3.0x3.0. run 4429
  • 172868 05/04/06 04:49 S. Frullani NMR measurement: Pol = 55.0 %
  • 172867 05/04/06 04:43 Armando Acha MWDC Currents. Run 4428
  • 172866 05/04/06 04:42 A. Acha Happex FM. Run 4428
  • 172865 05/04/06 04:41 Armando Acha Raster size 2.0x2.0 . 3He
  • 172861 05/04/06 03:10 Armando Acha MWDC Currents . Run 4427
  • 172859 05/04/06 03:08 A. Acha Raster 2x2.Run 4427
  • 172858 05/04/06 03:02 A. Kolarkar Analysis Shift Summary 2300 - 0300
  • 172857 05/04/06 03:00 Armando Acha Run 4426
  • 172854 05/04/06 02:06 Armando Acha MWDC Currents Run 4426
  • 172853 05/04/06 02:04 Armando Acha Raster size 2x2 Run 4426
  • 172849 05/04/06 01:46 Armando Acha MWDC Currents. RUN 4425
  • 172848 05/04/06 01:45 Armando Acha run 4425. raster size
  • 172844 05/04/06 01:38 Armando Acha Beam Current Monitor
  • 172843 05/04/06 01:37 Armando Acha Happex FM for run 4424
  • 172842 05/04/06 01:34 A. Acha Correction for entries 172840 and 172841
  • 172841 05/04/06 01:31 Armando Acha MWDC Currents .Run 4424
  • 172840 05/04/06 01:30 Armando Acha run 4424. raster 3x3. H2
  • 172836 05/04/06 01:14 Armando Acha MWDC run 4423. Target H2
  • 172835 05/04/06 01:13 A. Acha target H2 run 4423 raster 3.0x3.0
  • 172831 05/04/06 00:54 A. Acha MWDC Currents
  • 172830 05/04/06 00:53 A. Acha raster 3.0x3.0 run 4422
  • 172826 05/04/06 00:43 A. Acha MWDC Currents run 4421
  • 172825 05/04/06 00:39 A. Acha raster run 4421
  • 172821 05/04/06 00:29 Nilanga MWDC currents for 2 x 2 raster
  • 172820 05/04/06 00:28 Nilanga MWDC currents for 1.5 x 1.5 raster
  • 172819 05/04/06 00:28 Nilanga spot ++ for run 4420
  • 172815 05/04/06 00:07 A. Camsonne HV neutron detector for veto
  • 172814 05/04/06 00:06 Nilanga raster sizes
  • Wednesday

  • 172812 05/03/06 23:59 frw Swing Shift Summary
  • 172811 05/03/06 23:58 A. Camsonne Compton tune postponed to tomorrow
  • 172810 05/03/06 23:49 Nerses run 4418 raster
  • 172808 05/03/06 23:45 frw quick raster scan on reference cell tune
  • 172806 05/03/06 23:41 B. Craver target to reference cell position
  • 172799 05/03/06 23:29 frw raster scan summary
  • 172798 05/03/06 23:27 Nerses spot++ run 4416, raster 4x4
  • 172794 05/03/06 23:23 A. Camsonne Compton background after beam resumed
  • 172793 05/03/06 23:21 B. Craver MWDC currents - 3.5x4.5 mm
  • 172792 05/03/06 23:20 Nerses spot++ run 4415, raster size 4x4
  • 172789 05/03/06 23:17 B. Craver MWDC currents - 4x4 mm
  • 172787 05/03/06 23:16 Nerses spot++ run 4414
  • 172784 05/03/06 23:15 B. Craver MWDC currents on 3x3 mm
  • 172782 05/03/06 23:12 Nerses spot++ run 4413
  • 172781 05/03/06 23:12 B. Craver MWDC currents
  • 172775 05/03/06 23:02 B. Craver spot++
  • 172770 05/03/06 22:49 B. Craver MWDC currents
  • 172767 05/03/06 22:43 B. Craver spot++
  • 172762 05/03/06 22:29 frw start checkout
  • 172757 05/03/06 22:15 A. Camsonne HV neutron detector adjustment
  • 172750 05/03/06 21:39 rsholmes Beam update
  • 172749 05/03/06 21:36 B. Craver target in empty position
  • 172746 05/03/06 20:22 rsholmes Only slightly wrong
  • 172745 05/03/06 20:09 B. Craver Helicity issue
  • 172744 05/03/06 19:58 B. Craver Helicity problems remain
  • 172743 05/03/06 19:56 B. Craver target computer crashed
  • 172738 05/03/06 19:47 B. Craver restore thresholds
  • 172737 05/03/06 19:46 B. Craver Beam is back
  • 172734 05/03/06 19:10 B. Craver Helicity warnings
  • 172730 05/03/06 18:06 A. Camsonne Neutron detector HV adjustment
  • 172729 05/03/06 17:56 B. Craver beam permit
  • 172726 05/03/06 17:38 frw hall now in power permit
  • 172725 05/03/06 16:35 kelleher NMR measurement: Pol = 54.2 %
  • 172724 05/03/06 16:12 rsholmes Beam progress
  • 172714 05/03/06 10:41 Katich NMR measurement: Pol = 56.3 %
  • 172700 05/03/06 01:41 J. Miller Neutron Database
  • Tuesday

  • 172692 05/02/06 17:53 frw Swing Shift Summary
  • 172691 05/02/06 17:51 frw no beam until 7am
  • 172689 05/02/06 17:49 B. Craver Cosmics settings
  • 172688 05/02/06 17:46 frw controlled access
  • 172687 05/02/06 17:46 frw BB DC HV on
  • 172686 05/02/06 17:34 B. Craver target computer crash
  • 172685 05/02/06 16:29 frw BB DC HV off
  • 172681 05/02/06 15:58 Saha Day Shift Summary
  • 172680 05/02/06 15:23 B. Craver NMR measurement: Pol = 53.2 %
  • 172676 05/02/06 12:33 riordan Momentum Raster Dependence
  • 172674 05/02/06 08:26 Saha Hall A in Controlled Access
  • 172673 05/02/06 08:25 A. Camsonne Filters on additionnal channels for ND and pedestal file update for neutron detector
  • 172672 05/02/06 08:17 A. Kolarkar NMR measurement: Pol = 49.1 %
  • 172660 05/02/06 00:50 E.Chudakov MPS signal fan-out is half-broken
  • Monday

  • 172659 05/01/06 22:25 Samuel target computer crashed
  • 172658 05/01/06 22:22 A. Camsonne Compton Led runs Amplitude 1.82V
  • 172654 05/01/06 22:16 frw Swing Shift Summary
  • 172652 05/01/06 22:14 frw now in controlled access
  • 172651 05/01/06 21:43 frw no beam tonite
  • 172650 05/01/06 21:11 A. Camsonne Compton LED run 13605/1352
  • 172649 05/01/06 21:01 A. Camsonne Compton led run 100 KHz
  • 172648 05/01/06 20:41 A. Camsonne Compton run led 5 KHz and 25 KHz
  • 172647 05/01/06 19:30 A. Camsonne Compton LED runs
  • 172646 05/01/06 17:59 kelleher turned lasers on
  • 172645 05/01/06 17:56 A. Camsonne Integrating Compton 5 KHz LED test run 1338
  • 172644 05/01/06 17:34 kelleher Helium spins flipped
  • 172643 05/01/06 15:58 O. Glamazdin shift summary, day
  • 172641 05/01/06 14:17 R. Feuerbach correction: det1 was connected (for pulsed beam studies)
  • 172640 05/01/06 14:11 A. Kolarkar Target Lasers Helicity Flipped
  • 172639 05/01/06 14:04 A. Kolarkar Target Computer Crash Update - solution applied
  • 172636 05/01/06 12:55 A. Kolarkar Target Computer Crash and Microsoft®'s Online Error Analysis
  • 172632 05/01/06 12:17 J. Miller Pedestal Run for BH
  • 172628 05/01/06 11:55 S.Abrahamyan trigger for kinematics 1b
  • 172624 05/01/06 08:30 kelleher evacuated reference cell
  • 172623 05/01/06 08:28 A. Camsonne Analysis shift summary 6-8 AM
  • 172622 05/01/06 07:59 Aniol shift summary, owl
  • 172616 05/01/06 07:15 Aniol beam off for restricted access
  • 172615 05/01/06 07:15 A. Camsonne Threshold set for cosmics data taking
  • 172609 05/01/06 06:27 Aniol beam off/acc magnet
  • 172608 05/01/06 06:27 Aniol change os1 to 10^5 after 4378
  • 172607 05/01/06 06:10 J. Miller Analysis Shift Summary 2 am - 6 am
  • 172603 05/01/06 05:07 Aniol beam back
  • 172602 05/01/06 05:04 Aniol beam off- accelerator magnet
  • 172601 05/01/06 04:33 Aniol why was ps1 changed to 10^6?
  • 172600 05/01/06 04:15 S. Frullani NMR measurement: Pol = 53.5 %
  • 172596 05/01/06 03:37 A. Kolarkar Analysis Shift Summary - 0300
  • 172586 05/01/06 00:34 Aniol spot++
  • 172584 05/01/06 00:17 Nerses Shift summary 16:00-24:00